Shift summary eve 7/16/99

Julie Whitmore (jaws@fnpxsr)
Sat, 17 Jul 1999 15:43:11 CDT

Hi,
Apparently some people did not receive this. -julie

------- Forwarded Message

Date: Sat, 17 Jul 1999 01:19:08 -0400 (EDT)
From: Lisa Manning <manning@phys.virginia.edu>
Subject: Shift summary eve 7/16/99
To: shiftsum@vivsgi
Message-id: <Pine.OSF.3.95.990717011748.24823A-100000@uvahea.phys.virginia.edu>
MIME-version: 1.0
Content-type: TEXT/PLAIN; charset=US-ASCII

Shift Summary for eve shift, 7/16/99

On Shift: Julie, Lisa Manning

Hours of beam available at all intensities: 8.25
Hours of beam available at nominal intensity: 8.25
Hours of E832 data (special runs not included): 7.75
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time:
DLT drive swap and Booster vacuum problems

========================================================================
Runs ended this shift: none

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:none

========================================================================
Overview of shift: Smooth e832 running.

========================================================================
Description of Shift:

run spill time comment
- ------------------------------------------------------------------------
13924 15:40 Start Beam 832 run

5-12 16:05 trigger stop. 'AZTEC busy'
Multiple kfends fail to cure
problem. Clear busy on crate 53 by
unplugging/reseating controller cable.

~38 16:45 Beam goes away and comes back 2 spills later

17:25 Rick K and Julie change the blue
ice on the alcohol bubbler and
discover that the temperature
immediately changes from -5.0 to
-3.9 when the plastic housing surrounding
the bubbler is touched.
Apparently, the platic housing is not seated
properly and leans against the 'set point'
toggle button, causing the bubbler to read
the set point temperature(-5.0) instead of
the actual temperature(-3.9).DC experts
notified.

135-39 19:00 CSI channel 950 drops out of read-out.
Pipeline initialization and kfrend fix
the problem. 4 spills lost.

~135 19:00 Notice boa reads 2.2-2.6,but returns to
normal|(~1.39) after CSI drama.

19:45 Notice cosmic T16 and T17 have low ADC value
and low number of hits per counter. Also, adc
pedestal apparently corresponding to T17 is
shifted ~35 counts

19:50 New bottle of CO2 online

20:10 Marty from MCR dropped by on his 'Listening
Tour'. Nice chat about beam, Julie gets to
discuss nuances of 97 run beam

21:50 Julie changes blue ice again. Alcohol temp is
-3.3. Gas shed temperature is 86 F, compared
to 90 F before gas shed door was opened
(Outside temp a balmy 80 F)

23:15 Modified e832_shift_check list to include in
online histogram kumacs to run regmon#muonrun.

336 23:45 Channel 13. Booster vacuum problems.
Estimate 15 mins.

360 00:00 Beam is back

00:00 Hand over to owl shift

========================================================================
Follow up Notes on Previous Shift
HCC did not cause any problems
ArEth replaced

========================================================================
Items to Follow up on Next Shift
Keep an eye on the alcohol bubbler temperature. Make sure that
the plastic housing is not leaning on the toggle button for the set
point.

------- End of Forwarded Message